### v1

# Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction

Abstract
Aspect sentiment quadruple prediction (ASQP) facilitates a detailed understanding of opinions expressed in a text by identifying the opinion term, aspect term, aspect category and sentiment polarity for each opinion. However, annotating a full set of training examples to fine-tune models for ASQP is a resource-intensive process. In this study, we explore the capabilities of large language models (LLMs) for zero- and few-shot learning on the ASQP task across five diverse datasets. We report F1 scores slightly below those obtained with state-of-the-art fine-tuned models but exceeding previously reported zero- and few-shot performance. In the 40-shot setting on the Rest16 restaurant domain dataset, LLMs achieved an F1 score of 52.46, compared to 60.39 by the best-performing fine-tuned method MVP. Additionally, we report the performance of LLMs in target aspect sentiment detection (TASD), where the F1 scores were also close to fine-tuned models, achieving 66.03 on Rest16 in the 40-shot setting, compared to 72.76 with MVP. While human annotators remain essential for achieving optimal performance, LLMs can reduce the need for extensive manual annotation in ASQP tasks.
Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction
Nils Constantin Hellwig Media Informatics Group University of Regensburg Regensburg, Germany nils-constantin.hellwig@ur.de                        Jakob Fehle Media Informatics Group University of Regensburg Regensburg, Germany jakob.fehle@ur.de
Udo Kruschwitz Information Science Group University of Regensburg Regensburg, Germany udo.kruschwitz@ur.de                        Christian Wolff Media Informatics Group University of Regensburg Regensburg, Germany christian.wolff@ur.de
1 Introduction
Transformer-based large language models (LLMs) have gained significant attention due to their capability to address a broad spectrum of natural language processing (NLP) tasks, such as text summarization, translation, reading comprehension and text classification (Brown et al., 2020; Dubey et al., 2024). Noteworthy LLMs include Llama-3.1 (Dubey et al., 2024), Gemma-2 (Gemma et al., 2024), and Mixtral (Jiang et al., 2024), which are accessible in various parameter sizes with open model weights and commercial models like GPT-4 (Achiam et al., 2023) and Claude 3 (Anthropic, 2024).
Previous research explored zero- and few-shot scenarios in which the LLM generates outputs with either none or only a few labelled examples provided in the prompt Gou et al. (2023); Zhang et al. (2024). This eliminates the need for supervised model training, such as for small language models [^1] There is no universally accepted definition for categorizing language models as small or large. As handled by Zhang et al. (2024), models with fewer than 1 billion parameters are considered small, while those with 1 billion or more parameters are classified as large. (SLMs) using annotated datasets (Wang et al., 2023c). This approach is particularly appealing because data annotation is often deemed complex and expensive, both in terms of time or financial cost, thereby complicating the development of text classification solutions tailored to specific tasks (Fehle et al., 2023; Gretz et al., 2023; Li et al., 2023).
An extensively studied task in NLP where manual annotations pose significant challenges is aspect-based sentiment analysis (ABSA) (Zhang et al., 2022). This task facilitates the understanding of customer opinions expressed in reviews or feedback (Pontiki et al., 2014). Unlike traditional sentiment classification, which assigns a single sentiment label (commonly positive, negative, or neutral) to an entire text document, ABSA requires annotators to identify all aspects within the text and determine the sentiment associated with each one (Zhang et al., 2022).
A prominent subtask of ABSA is aspect sentiment quad prediction (ASQP), which provides exceptionally detailed insights into the author’s opinions by identifying four sentiment elements for each opinion: aspect term (a), aspect category (c), sentiment polarity (p) and opinion term (o) (Zhang et al., 2021a). Consequently, the annotation process for training examples is highly demanding, particularly when multiple opinions need to be annotated within a single text.
Previous research has predominantly concentrated on 0- to 10-shot learning, exclusively utilizing the English-language restaurant domain datasets Rest15 and Rest16 introduced by Zhang et al. (2021a).
In this study, we extend the analysis to include up to 50 few-shot examples and evaluate the approach on a diverse series of five datasets. The datasets utilized in this work include Rest15 and Rest16, introduced by Zhang et al. (2021b) and we incorporate the OATS dataset by Chebolu et al. (2024), which consists of hotel reviews from TripAdvisor and online learning reviews collected from Coursera. Finally, we introduce a novel ASQP dataset, comprising annotated sentences from airline reviews, which is published as part of this work.
We considered the following research questions:
RQ1: How does varying the number of few-shot examples (from 0 to 50) impact performance on the ASQP task?
RQ2: How do LLMs perform on the ASQP task compared to SLMs trained on annotated examples?
RQ3: Does self-consistency (SC) prompting (Wang et al., 2022a), where multiple outputs are generated from the same prompt and the most consistent response is selected, improve performance on the ASQP task?
We employed Google’s Gemma-2-27B (Gemma et al., 2024) and report the performance for the smaller-sized Gemma-2-9B. In addition, we report the LLMs’ performance on the target aspect sentiment detection (TASD) task, which focuses on the identification of (a, c, p)-triplets. All code and results of this study is publicly available on GitHub [^2] https://anonymous.4open.science/r/llm-absa-quad-zero.
2 Related Work
Strategy Method ASQP TASD
Rest15 Rest16 Rest15 Rest16
Zero-shot learning gpt-3.5-turbo, 0-shot (uncased) (Gou et al., 2023) 22.87 - - 34.08
gpt-3.5-turbo, 0-shot (Zhang et al., 2024) 10.46 14.02 - -
text-davinci-003, 0-shot (Zhang et al., 2024) 13.73 18.18 - -
ChatABSA, 0-shot (Bai et al., 2024) 27.11 30.42 39.21 41.28
Few-shot learning gpt-3.5-turbo, 1-shot (Zhang et al., 2024) 30.15 31.98 - -
gpt-3.5-turbo, 5-shot (Zhang et al., 2024) 31.21 38.01 - -
gpt-3.5-turbo, 10-shot (uncased) (Gou et al., 2023) 34.27 - - 46.51
gpt-3.5-turbo, 10-shot (Zhang et al., 2024) 30.92 40.15 - -
ChatABSA, 1-shot (Bai et al., 2024) 28.13 33.84 37.23 41.92
ChatABSA, 5-shot (Bai et al., 2024) 33.26 31.92 43.00 45.04
ChatABSA, 10-shot (Bai et al., 2024) 32.14 33.26 45.93 47.00
Fine-tuning TAS-BERT (Wan et al., 2020) 34.78 43.71 57.51 65.89
Extract-Classify (Cai et al., 2021) 36.42 43.77 - -
GAS (Zhang et al., 2021b) 45.98 56.04 60.63 68.31
Paraphrase (Zhang et al., 2021a) 46.93 57.93 63.06 71.97
DLO (Hu et al., 2022) 48.18 59.79 62.95 71.79
MVP (Gou et al., 2023) 51.04 60.39 64.53 72.76
Table 1: Performance on the ASQP and TASD task. F1 scores of both LLM-based and fine-tuned approaches from related work.
2.1 Aspect Sentiment Quad Prediction
Figure 1: Annotated example for ASQP from Rest16 (Zhang et al., 2021a). One or multiple opinion-quadruple annotations are assigned to each sentence.
The development of methodologies for addressing the ASQP task was strongly influenced by the work of Zhang et al. (2021a), which introduced two annotated datasets for the ASQP task, Rest15 and Rest16. An example of such annotations is illustrated in Figure 1. Both datasets comprise annotated sentences derived from restaurant reviews. The annotations are sourced from the SemEval Shared Task datasets from 2015 and 2016 (Pontiki et al., 2015, 2016), which originally included only (a, c, p)-triplets and thus did not include annotations for opinion terms.
Since the release of Rest15 and Rest16, generative methods within a unified framework have emerged as the state-of-the-art (SOTA) approach for the ASQP task. Various strategies have been explored to generate sentiment elements in specific formats that exploit label semantics. These include approaches employing structured extraction schemas (Lu et al., 2022), sequential representations of sentiment elements (Gou et al., 2023) and natural language formats (Gou et al., 2023; Liu et al., 2021), wherein quadruples are systematically converted into natural language sentences. Performance scores for these methods are presented in Table 1.
All the aforementioned approaches rely on small text generation models, such as t5-base (Raffel et al., 2020), which utilizes an encoder-decoder architecture based on the transformer architecture Vaswani (2017). The t5-base model, comprising 223 million parameters, is fine-tuned specifically for the ASQP task.
2.2 Large Language Models for Aspect-based Sentiment Analysis
The zero- and few-shot capabilities of LLMs have been demonstrated across various NLP tasks, e.g. question answering (Chada and Natarajan, 2021; Brown et al., 2020), named entity recognition (Cheng et al., 2024; Wang et al., 2023b), information retrieval (Faggioli et al., 2023; Wang et al., 2022b) or sentiment analysis (Zhang et al., 2024). In many cases, these models have achieved performance scores comparable to fine-tuned approaches, with few-shot learning often outperforming zero-shot learning.
In the domain of ABSA, LLMs have been employed in both zero- and few-shot settings. However, these efforts were constrained to a maximum of 10 few-shot examples within the prompt’s context, addressing both ASQP and ABSA tasks with fewer sentiment elements (Conneau, 2019; Gou et al., 2023; Zhang et al., 2024).
Zhang et al. (2024) employed OpenAI’s gpt-3.5-turbo (Brown et al., 2020) for End-to-End ABSA (E2E-ABSA, focus on (a, p) pairs) and achieved an F1 score of 54.46 and 63.30 on the Rest14 dataset (restaurant domain) from Pontiki et al. (2014) for zero- and 10-shot learning, respectively. A fine-tuned t5-large model (Raffel et al., 2020) achieved a slightly higher F1 score of 75.31. Similarly, Wu et al. (2024) analysed multiple open source LLMs with less than 10 billion parameters, as well as commercial LLMs for multilingual E2E-ABSA in a zero-shot setting. In multilingual ABSA, applying prompting strategies such as chain-of-thought (CoT) prompting did not improve performance when averaged across the LLMs considered. However, the best performing LLM, GPT-4o-CoT, achieved an F1 score of 52.81 which is slightly below the performance of the most performant fine-tuned model XLM-R (Conneau, 2019) (68.86). Wu et al. (2024) also evaluated a self-consistency (SC) prompting strategy, where the most frequent label across five generated outputs was selected as the final label. SC did not lead to an improvement in the performance.
With regard to the ASQP task, Zhang et al. (2024) achieved F1 scores below 20 for both Rest15 and Rest16 (see Table 1). Performance was improved to F1 scores above 30 on both Rest15 and Rest16 by providing 1, 5 or 10 few-shot examples.
Gou et al. (2023) surpassed the performance reported by Zhang et al. (2024) and reported an F1 score of 22.87 (zero-shot) and 34.27 (10-shot) on the Rest16 dataset, slightly exceeding the performance reported by Zhang et al. (2024). Notably, Gou et al. (2023) presented the sentences to be annotated and few-shot examples in an uncased format within the prompt, differing from the approach by Zhang et al. (2024). Furthermore, the task descriptions were formatted differently, with Gou et al. (2023) offering descriptions on each of the four sentiment elements considered in the respective ABSA task. Bai et al. (2024) adopted a distinct approach (referred to as ChatABSA) to processing its outputs, leading to performance improvements in the zero-shot setting but not in the few-shot setting. In the prompt, it was stated that the output should be in the JSON format. Furthermore, predicted aspect terms or opinion terms that were not explicitly mentioned in the original sentence were systematically set to null.
In summary, previous studies demonstrated that few-shot learning massively boosts performance in ABSA tasks but did not exceed the performance of models fine-tuned on annotated examples.
3 Methodology
We utilized LLMs to tackle the ASQP task across 0-, 10-, 20-, 30-, 40-, and 50-shot settings on different datasets. The performance is compared to that achieved using a dedicated training set to fine-tune smaller pre-trained language models. Furthermore, we report performance results for the TASD task.
3.1 Evaluation
3.1.1 Datasets
Rest15 Rest16 FlightABSA OATS Coursera OATS Hotels
# Train 834 1,264 1,351 1,400 1,400
# Test 537 544 387 400 400
# Dev 209 316 192 200 200
# Aspect Categories 13 13 13 28 33
Language en en en en en
Domain restaurant restaurant airline e-learning hotel
Table 2: Overview of all ASQP datasets considered for evaluation. The datasets cover a range of different numbers of considered aspect categories and domains.
Table 2 presents an overview of the datasets used in this study, including Rest15 and Rest16, along with three additional datasets covering diverse domains.
Rest15 & Rest16: ASQP annotations originate from Zhang et al. (2021a) and the TASD annotations from Wan et al. (2020). This ensured comparability with the performance scores reported in previous research.
FlightABSA: A novel dataset containing 1,930 sentences annotated for ASQP. Properties of the annotated dataset are provided in Appendix B.
OATS Hotels & OATS Coursera: We utilized a subset of two corpora recently introduced by Chebolu et al. (2024) comprising ASQP-annotated sentences from reviews on hotels and e-learning courses. A detailed description of the data preprocessing for the OATS datasets can be found in Appendix A.
For the TASD task, we removed the opinion terms from the quadruples in annotations from FlightABSA, OATS Coursera and OATS Hotels. Subsequently, any duplicate triplets (a, c, p) that appeared twice in a sentence were discarded.
3.1.2 Setting
For evaluation, the test dataset was considered for all datasets. An LLM was prompted five times with different seeds (0 to 4) for each combination of ABSA task (ASQP and TASD), dataset and amount of random few-shot examples (0, 10, 20, 30, 40 or 50) taken from the training set in order to get five label predictions. For all seeds, the same few-shot examples were used; however, they were shuffled differently for each prompt execution. The average performance across all five runs is reported.
3.1.3 Metrics
As in previous works in the field of ABSA, we report the micro-averaged F1 score as well as precision and recall to assess the model’s performance. The F1 score is the harmonic mean of precision and recall. Precision measures the proportion of correctly predicted positive instances out of all instances predicted as positive (Jurafsky and Martin, 2024, p. 67). Recall quantifies the proportion of correctly predicted positive instances out of all actual positive instances in the dataset (Jurafsky and Martin, 2024, p. 67).
Similar to Zhang et al. (2021a), a quad prediction was considered correct if all the predicted sentiment elements are exactly the same as the gold labels. Recognizing the potential interest in class-level performance metrics for subsequent research, we have shared the predicted labels for every evaluated setting in our GitHub repository, allowing detailed class-level analysis.
Figure 2: The prompt includes both a task description and specification of the output format. The LLM is run with five different seeds and in the case of self-consistency prompting, the tuple that appears most often across the five predictions is incorporated into the final label.
3.2 Large Language Models
We employed Gemma-2-27B333google/gemma-2-27b: https://ollama.com/library/gemma2:27b by Google, which comprises 27.2 billion parameters (Gemma et al., 2024). Ollama444ollama: https://ollama.com was employed for inference, and the LLMs were loaded with 4-bit quantization. The model was chosen for its efficiency in terms of generated tokens per second, which is a critical factor given the extensive prompt execution requirements. Notably, our study required over 342,720 prompts to be executed, with many few-shot learning prompts encompassing over a thousand tokens. For larger models, such as Llama-3.3-70B Dubey et al. (2024), the required computational costs would have been hardly feasible with our resources. For comparison purposes, we also report performance for the smaller-sized LLM, Gemma-2-9B555google/gemma-2-9b: https://ollama.com/library/gemma2:9b.
The experiments were conducted on two NVIDIA RTX A5000 GPUs equipped with 24 GB of VRAM each. The LLM’s temperature parameter was set to 0.8 and generation was terminated upon encountering the closing square bracket character ("]") signifying the ending of a predicted label.
3.3 Prompt
3.3.1 Components
We adopted the prompting framework introduced by Gou et al. (2023) with some modifications. The employed prompt is illustrated in Figure 2 and an example is provided in Appendix C. The main components of the prompt include a list of explanation on all considered sentiment elements and the specification of the output format.
Unlike Gou et al. (2023), our prompt instructs the LLM to pay attention to case sensitivity when returning aspect and opinion terms. Hence, the identified phrases should appear in the predicted tuple as they do in the sentence, similar to all supervised approaches mentioned in the related work section. Therefore, in the prompt, we clearly stated that the exact phrases should appear in the predicted label.
Since we execute each prompt with five different seeds, we also report the performance when employing the self-consistency prompting technique introduced by Wang et al. (2022a). The key idea is to select the most consistent answer from multiple prompt executions. We adapted the approach for ABSA by incorporating a tuple into the merged label if it appears in the majority of the predicted labels. As illustrated in Figure 2, this corresponds to a tuple appearing in at least 3 out of 5 predicted labels.
3.4 Output Validation
Since LLMs such as Gemma-2-27B cannot be strictly constrained to a fixed output format, we programmatically validated the output of the LLM. For the predicted label, several criteria needed to be met for the generation to be considered valid:
  • Format: The output must be a list of one or more tuples consisting of strings (quadruples for ASQP, triplets for TASD).
  • Sentiment: The sentiment must be either ’positive’, ’negative’ or ’neutral’.
  • Aspect category: Only the categories considered for the respective dataset and thus being mentioned in the prompt should be predicted as a part of a tuple.
  • Aspect and opinion terms: Both must appear in the given sentence as predicted.
If any of the specified criteria for reasoning or label validation is not met, a regeneration attempt was triggered. If the predicted label was still invalid after 10 attempts, an empty label ([]) was considered as the predicted label.
3.5 Baseline Model
We compared the previously mentioned zero- and few-shot conditions against three SOTA baseline approaches, which are, the three best-performing methods for ASQP and TASD on the Rest15 and Rest16 datasets: Paraphrase (Zhang et al., 2021a), DLO (Hu et al., 2022) and MVP (Gou et al., 2023).
Paraphrase (Zhang et al., 2021a):
Paraphrase is used to linearize sentiment quads into a natural language sequence to construct the input target pair.
DLO (Hu et al., 2022):
Dataset-level order is a method designed for ASQP that leverages the order-free property of quadruplets. It identifies and utilizes optimal template orders through entropy minimization and combines multiple effective templates for data augmentation.
MVP (Gou et al., 2023):
Multi-view-Prompting introduces element order prompts. The language model is guided to generate multiple sentiment tuples, with a different element order each, and then selects the most reasonable tuples by a voting mechanism. This method is highly resource-intensive, as multiple input-output pairs are created for each example in the train set, each comprising different sentiment element positions.
For all three approaches, we conducted training using the entire dataset and performed training with only 10, 20, 30, 40, or 50 training examples equally to the ones employed for the few-shot learning conditions. Training was conducted using five different random seeds (0 to 4). Moreover, to facilitate comparisons across datasets, we trained models using 800 training examples, as this represents the largest multiple of 100 examples available for all train sets (900 training examples are not available for Rest15). The results obtained using the full training sets of Rest15 and Rest16 were extracted from the works of Zhang et al. (2021a), Hu et al. (2022), and Gou et al. (2023).
For all methods, we used the hyperparameter configurations used by Zhang et al. (2021a), Hu et al. (2022) and Gou et al. (2023). The only exception was the 10-shot condition, where batch size was set to 8 instead of 16, as the limited number of examples (10) could not form a batch of 16 examples.
Method Prompting Strategy # Few-Shot / # Train Rest15 Rest16 FlightABSA
OATS
Coursera
OATS
Hotels
F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec
Gemma-2-9B - 0 13.26 14.11 12.50 16.94 17.86 16.12 33.41 35.68 31.42 9.58 9.52 9.64 21.06 24.76 18.34
10 20.09 18.78 21.61 20.61 19.34 22.05 23.84 24.32 23.39 5.78 5.43 6.18 17.95 20.97 15.70
20 23.85 23.31 24.43 27.22 26.41 28.09 24.05 23.94 24.17 8.80 8.49 9.12 21.58 22.65 20.61
30 25.62 25.12 26.14 31.66 31.00 32.37 29.87 29.74 30.00 11.65 11.29 12.03 23.69 25.33 22.25
40 26.31 25.73 26.92 30.07 29.69 30.46 31.56 31.40 31.73 12.28 12.21 12.35 25.13 27.43 23.19
50 27.32 27.01 27.65 31.89 31.43 32.37 33.77 33.14 34.44 11.73 11.51 11.95 27.23 30.11 24.85
SC 0 14.58 17.12 12.70 16.82 19.76 14.64 36.86 43.08 32.20 11.24 13.64 9.56 21.74 28.77 17.48
10 26.84 49.83 18.36 25.63 49.64 17.27 30.12 62.57 19.83 3.29 20.00 1.79 21.75 47.00 14.15
20 30.56 56.04 21.01 34.76 57.51 24.91 31.58 60.58 21.36 10.82 43.66 6.18 26.92 58.60 17.48
30 33.74 55.98 24.15 39.53 59.75 29.54 38.73 62.98 27.97 18.78 54.29 11.35 30.75 60.08 20.67
40 35.48 57.46 25.66 37.79 56.93 28.29 39.54 60.70 29.32 15.75 49.47 9.36 30.61 57.92 20.80
50 34.19 57.06 24.40 40.56 59.90 30.66 40.68 61.72 30.34 16.34 45.45 9.96 31.65 57.93 21.78
Gemma-2-27B - 0 27.74 26.80 28.75 35.79 34.80 36.85 37.09 37.13 37.05 11.74 10.71 12.99 27.51 28.62 26.49
10 35.48 33.38 37.86 37.73 35.71 40.00 40.74 40.18 41.32 18.08 17.05 19.24 28.78 30.45 27.30
20 33.01 32.14 33.94 44.01 42.13 46.08 38.67 38.04 39.32 20.03 19.15 21.00 31.95 32.62 31.32
30 35.21 34.12 36.38 44.92 43.22 46.76 43.54 42.59 44.54 21.14 20.28 22.07 36.05 36.69 35.42
40 36.19 34.69 37.84 46.71 45.22 48.31 42.01 41.21 42.85 21.40 20.63 22.23 34.26 34.85 33.70
50 36.31 34.97 37.76 46.14 45.12 47.21 44.05 42.70 45.49 21.23 20.42 22.11 35.05 36.58 33.65
SC 0 29.74 30.74 28.81 38.60 40.14 37.17 38.96 41.21 36.95 13.40 13.89 12.95 29.27 32.49 26.63
10 39.34 41.69 37.23 41.11 43.72 38.80 44.44 50.00 40.00 22.81 28.70 18.92 30.92 38.59 25.80
20 36.35 40.72 32.83 48.48 51.33 45.93 43.21 48.72 38.81 25.67 33.23 20.92 38.84 48.06 32.59
30 40.33 45.10 36.48 49.73 53.21 46.68 49.15 55.56 44.07 26.24 34.64 21.12 40.75 49.41 34.67
40 41.50 46.49 37.48 52.46 55.87 49.44 47.13 53.08 42.37 25.22 34.98 19.72 39.64 49.28 33.15
50 40.62 46.66 35.97 51.24 56.97 46.56 49.91 56.14 44.92 27.20 36.99 21.51 40.24 52.57 32.59
MVP (Gou et al., 2023) 10 10.58 12.00 9.46 12.37 14.40 10.84 9.38 11.66 7.84 12.88 14.46 11.62 6.98 8.42 5.97
20 18.71 21.22 16.73 21.49 24.30 19.27 14.27 17.43 12.09 18.85 20.79 17.25 14.30 16.03 12.92
30 24.36 26.54 22.52 27.58 30.83 24.96 22.53 26.82 19.42 21.32 23.25 19.68 20.89 23.17 19.03
40 25.95 27.72 24.40 32.72 33.56 31.94 28.15 32.17 25.03 20.21 22.02 18.68 24.71 27.00 22.78
50 30.20 31.07 29.38 33.32 34.75 32.02 33.12 35.09 31.38 22.07 24.16 20.32 29.91 33.08 27.31
800 50.02 48.99 51.09 58.09 56.31 59.97 57.46 56.23 58.74 30.26 29.91 30.62 53.37 52.41 54.36
Full 51.04 - - 60.39 - - 57.90 56.09 59.83 32.50 32.04 32.97 55.03 54.38 55.69
DLO (Hu et al., 2022) 10 1.32 1.64 1.11 3.56 4.02 3.23 3.44 4.34 2.85 4.75 5.35 4.26 2.63 3.66 2.06
20 5.48 6.78 4.60 11.14 10.54 11.91 3.48 4.39 2.88 9.51 10.64 8.61 5.34 6.36 4.65
30 9.47 9.54 9.46 7.18 8.44 6.28 3.60 4.55 2.98 11.39 12.84 10.24 5.13 6.48 4.26
40 17.61 17.07 18.19 20.15 20.69 19.67 13.81 15.09 12.78 16.43 17.79 15.26 14.96 15.99 14.08
50 25.55 24.58 26.62 23.50 23.75 23.25 17.98 18.58 17.42 19.38 20.72 18.21 23.09 23.67 22.59
800 46.32 45.61 47.07 56.88 55.65 58.17 54.96 54.10 55.86 30.79 30.63 30.96 53.65 52.57 54.77
Full 46.93 46.16 47.72 57.93 56.63 59.30 57.76 57.37 58.17 32.34 32.06 32.63 53.87 52.61 55.19
Paraphrase (Zhang et al., 2021a) 10 3.88 4.14 3.65 4.49 4.79 4.23 4.46 5.63 3.69 5.88 6.62 5.30 3.60 3.89 3.36
20 12.38 14.80 10.64 14.01 14.56 13.52 9.67 12.15 8.03 10.28 11.16 9.52 7.53 6.47 9.10
30 18.94 18.34 19.62 23.62 24.37 22.93 17.43 18.98 16.14 16.92 17.59 16.29 17.80 17.45 18.17
40 22.93 21.88 24.10 26.72 25.73 27.78 24.31 26.76 22.27 16.99 18.04 16.06 23.29 23.95 22.69
50 27.93 25.94 30.26 29.63 29.23 30.06 28.81 28.14 29.53 19.87 21.64 18.37 28.14 29.46 26.93
800 48.76 47.54 50.04 58.59 56.65 60.68 56.91 55.33 58.58 29.96 29.54 30.40 53.97 53.23 54.73
Full 48.18 47.08 49.33 59.79 57.92 61.80 58.61 56.85 60.47 32.01 31.48 32.55 55.62 54.45 56.84
Table 3: Performance scores for ASQP. For the Rest15 and Rest16 datasets, performance scores achieved when employing the full training set ("Full") are taken from Gou et al. (2023), Hu et al. (2022) and Zhang et al. (2021b) for MVP, DLO and Paraphrase, respectively. The best score achieved by a method is presented in bold.
4 Results
The performance scores for the evaluated configurations are shown in Table 3 for the ASQP task and in Appendix D for the TASD task. Detailed performance scores focusing on individual sentiment elements are provided in Appendix E. Notably, for both tasks, we performed t-tests with Bonferroni correction (padj < .05) to examine whether significant differences exist between the F1 scores of the evaluated conditions (corresponding to the number of rows in Figure 3). No significant differences were observed.
Performance gains with an increasing number of few-shot examples. In most cases, increasing the number of few-shot examples resulted in incremental improvements in F1 scores across both ASQP and TASD tasks. The difference between zero- and few-shot prompting is substantial. For instance, on the Rest16 dataset under the SC prompting condition, the F1 score improved from 38.60 (0-shot) to 52.46 (40-shot). To further highlight this trend, we provide line plots (see Figure 4) that depict the influence of the number of few-shot examples on the F1 scores across all tasks, datasets, and models.
LLM performance slightly lower compared to SOTA fine-tuned approaches. For both TASD and ASQP, the performance achieved through zero- and few-shot prompting did not surpass that obtained when the entire training set was utilized. However, the best F1 scores achieved by Gemma-2-27B in the TASD task were often close to those achieved by fine-tuned approaches employing 800 or all examples from the training set. For example, on the Rest16 dataset, Gemma-2-27B achieved 66.03, which is slightly below the best F1 score achieved by a fine-tuned approach (MVP: 72.76). In case only 10 to 50 annotated examples were available for prompting or training, few-shot prompting consistently outperformed fine-tuning approaches across all sample sizes, with only a few exceptions.
Massive performance enhancements achieved through self-consistency. SC enabled considerable boosts of the F1 score, regardless of the amount of few-shot examples. However, recall was occasionally better without SC. Precision, on the other hand, was improved with SC in both tasks and across datasets. For instance, in the case of Gemma-2-9B, precision was increased by more than 20 percentage points in several instances.
The LLM’s parameter size matters. Gemma-2-9B demonstrated lower performance in terms of F1 scores for both ASQP and TASD. Across the five datasets, the F1 scores in the ASQP task were approximately 10 percentage points lower when using Gemma-2-9B instead of Gemma-2-27B. For example, on the Rest15 dataset, the best F1 score achieved with Gemma-2-9B was 35.48, while the best score for Gemma-2-27B was 41.50. Nonetheless, respectable F1 scores were still achieved with Gemma-2-9B, approaching to 50 (e.g., 49.96 for Rest15) and surpassing 50 (e.g., 54.77 for TASD on Rest16).
Lower performance in identifying opinion terms compared to other sentiment elements. As shown in the tables in Appendix E, performance in identifying sentiment (positive, negative, or neutral) is highly performant, with F1 scores exceeding 90. However, performance in identifying aspect and opinion terms is comparatively much lower.
5 Discussion
The results demonstrate performance improvements in F1 scores for both ASQP and TASD as the number of few-shot examples increases, highlighting the gap between zero- and few-shot prompting. in this chapter, we put the results of this work into the context of previous research and provide an outlook on the direction of future work.
New SOTA performance of LLMs. The performance scores reported in previous studies by Gou et al. (2023), Zhang et al. (2024) and Bai et al. (2024) for the ASQP task on the Rest15 and Rest16 datasets fall below those achieved by Gemma-2-27B in both zero- and 10-shot learning settings. The only exception is the TASD and Rest15, where ChatABSA (Bai et al., 2024) marginally outperforms Gemma-2-27B in zero-shot learning, with an F1 score of 39.21 compared to Gemma-2-27B’s F1 scores of 37.08 (with SC) and 35.98 (without SC). In contrast, on Rest16, Gemma-2-27B demonstrates a substantial advantage, achieving F1 scores of 53.13 (with SC) and 50.89 (without SC), exceeding ChatABSA’s F1 score of 41.28. Unlike prior studies, which have primarily evaluated up to 10-shot settings, we extended the investigation to a 10- to 50-shot setting for the first time. In this expanded range, Gemma-2-27B achieved notable F1 scores exceeding 50 for the ASQP task (e.g., Rest16 with SC: 52.46) and surpassing 60 for the TASD task (e.g., Rest16 with SC: 66.03). Notably, these substantial gains are primarily attributed to the use of SC prompting. Furthermore, this is in contrast to the work of Wu et al. (2024), whose SC approach for E2E-ABSA did not lead to an improvement in performance.
Model size and prompting strategy affect few-shot performance. Although Gemma-2-27B achieved competitive results in both ASQP and TASD, its performance remained slightly below fine-tuned SOTA approaches such as those by Gou et al. (2023), Zhang et al. (2021b), and Hu et al. (2022) when full training sets were employed. However, in scenarios with limited annotated examples, few-shot prompting consistently outperformed fine-tuning. The parameter size of the model also influenced performance, with Gemma-2-27B consistently outperforming its smaller counterpart, Gemma-2-9B.
Directions for enhancing low-resource task performance. Building on the promising results of this study, future research could focus on improving low-resource task performance through advanced prompt engineering techniques. Approaches such as chain-of-thought prompting (Wei et al., 2022) or plan-and-solve prompting (Wang et al., 2023a), which allowed for performance gains in other NLP tasks, hold significant potential. Furthermore, refining annotation guidelines or representing labels as natural language text, as proposed by Zhang et al. (2021b), could contribute to improved outcomes. Bigger LLMs, e.g. with 70B parameters, may provide additional performance benefits, given that our 27B model demonstrated superior results compared to the 9B variant.
Exploring less complex tasks and many-shot learning. In a broader context, future research could extend our approach to less complex tasks, in terms of the amount of considered sentiment elements, such as E2E-ABSA or aspect category sentiment analysis (ACSA) which focuses on aspect category and the sentiment expressed towards them. Beyond the low-resource setting considered in this study, one could explore the so-called "many-shot in-context learning" paradigm described by Agarwal et al. for ABSA, where hundreds or even the full training set is provided in the prompt. Observing that our approach achieved performance scores on the TASD task close to fine-tuned models, future work could investigate whether further increasing the number of shots lead to surpassing fine-tuned approaches.
Limitations
This study evaluated the performance of LLMs on ASQP tasks across a broad selection of datasets, few-shot settings, and LLM configurations. However, a limitation of this work is the selection of employed LLMs. We only employed LLMs comprising 9 or 27 billion parameters. Bigger-sized models such as Llama-3-70B (Dubey et al., 2024) or commercial models were not considered due to their prohibitive computational and financial costs. In order to evaluate each setting for a considered LLM, we executed a total of 171,360 prompts. Due to the amount of tokens in each prompt, the associated cost implications are substantial: about 191 hours (8 days) for Gemma-2-27B and 120 hours (5 days) for Gemma-2-9B. Hence, the time would further increase with an even bigger LLM in terms of parameter size. For commercial models such as GPT-4, executing all prompts would result in massive costs.
Finally, we must highlight the issue of potential data contamination, as it is the case for the previous studies introduced in the related work section. Meaning, it cannot be ruled out that the publicly available annotated datasets used in this study (except for FlightABSA) were included in the training data for both Gemma-2-9B and Gemma-2-27B.
Ethics Statement
All results and code used in this study are publicly available. The dataset we introduced, FlightABSA, is available upon request. We want to prevent the annotated dataset from being available online and then being inadvertently collected for pre-training LLMs.
Appendix A OATS Datasets Preprocessing
Since the OATS corpora, unlike Rest15 and Rest16 Zhang et al. (2021a), include examples where no quadruples were annotated, we excluded these instances as it is the case for Rest15, Rest16 and FlightABSA.
Two limitations of the OATS corpora led to a different approach for train-test-validation split. First, of the 7,188 (OATS Coursera) and 7,834 (OATS Hotels) training examples, 5,887 and 5,304 respectively included at least one annotated quadruple. Approaches relying on the training set would require significant training time when employing more than 5000 samples (which are also compared with the LLM’s performance scores). Secondly, the test set contained only 130 examples with at least one annotated quadruple. Due to these limitations, we decided to employ samples from the training set for our analysis. Hence, we took 2,000 examples from the training sets and a train-test-validation split (70:20:10) was applied.
Appendix B FlightABSA Dataset
Positive Negative Neutral Total
Aspect Category Explicit Implicit Explicit Implicit Explicit Implicit Explicit Implicit
AIRLINE#GENERAL 203 86 137 94 13 11 353 191
AIRLINE#PRICE 21 12 37 29 - - 58 41
AIRLINE#SERVICE 442 28 178 42 7 2 627 72
AIRPORT#OPERATION#BAGGAGE 14 - 38 1 1 - 53 1
AIRPORT#OPERATION#BOARDING 20 2 18 - - - 38 2
AIRPORT#OPERATION#CHECK_IN 50 - 25 1 4 - 79 1
ONBOARD#CLEANLINESS 24 1 11 - 1 - 36 1
ONBOARD#ENTERTAINMENT 19 1 13 - 4 - 36 1
ONBOARD#FOOD 68 - 67 1 15 1 150 2
ONBOARD#PRICE 5 - 5 1 - - 10 1
ONBOARD#SEAT#COMFORT 42 1 50 4 3 - 95 5
ONBOARD#SEAT#LEGROOM 32 - 21 - 2 - 55 -
PUNCTUALITY#GENERAL 42 18 95 51 1 1 138 70
Total 982 149 695 224 51 15 1728 388
((a)) Training set
Positive Negative Neutral Total
Aspect Category Explicit Implicit Explicit Implicit Explicit Implicit Explicit Implicit
AIRLINE#GENERAL 64 19 43 25 4 1 111 45
AIRLINE#PRICE 8 4 7 8 - - 15 12
AIRLINE#SERVICE 105 8 54 13 2 - 161 21
AIRPORT#OPERATION#BAGGAGE 3 - 13 - - - 16 -
AIRPORT#OPERATION#BOARDING 5 - 4 3 - - 9 3
AIRPORT#OPERATION#CHECK_IN 14 - 7 1 - - 21 1
ONBOARD#CLEANLINESS 6 1 2 - 1 - 9 1
ONBOARD#ENTERTAINMENT 4 - 7 - - - 11 -
ONBOARD#FOOD 21 - 26 - 4 - 51 -
ONBOARD#PRICE 1 - - 2 - - 1 2
ONBOARD#SEAT#COMFORT 10 - 16 3 1 - 27 3
ONBOARD#SEAT#LEGROOM 6 - 7 1 - - 13 1
PUNCTUALITY#GENERAL 16 4 18 17 1 - 35 21
Total 263 36 204 73 13 1 480 110
((b)) Test set
Positive Negative Neutral Total
Aspect Category Explicit Implicit Explicit Implicit Explicit Implicit Explicit Implicit
AIRLINE#GENERAL 32 16 18 15 1 1 51 32
AIRLINE#PRICE 3 3 4 4 - - 7 7
AIRLINE#SERVICE 61 5 30 7 2 - 93 12
AIRPORT#OPERATION#BAGGAGE 2 - 6 - - - 8 -
AIRPORT#OPERATION#BOARDING 5 - 2 - - - 7 -
AIRPORT#OPERATION#CHECK_IN 6 2 5 - - - 11 2
ONBOARD#CLEANLINESS 2 2 1 1 - - 3 3
ONBOARD#ENTERTAINMENT 2 - 2 - - - 4 -
ONBOARD#FOOD 13 - 7 - 1 - 21 -
ONBOARD#PRICE 1 - 2 - - - 3 -
ONBOARD#SEAT#COMFORT 5 - 12 2 - - 17 2
ONBOARD#SEAT#LEGROOM 5 - 2 - - - 7 -
PUNCTUALITY#GENERAL 4 1 17 6 - - 21 7
Total 141 29 108 35 4 1 253 65
((c)) Develop set
Table 4: Overview of FlightABSA. Aspect categories distribution per sentiment polarity and reference type.
B.1 Data Acquisition
FlightABSA comprises reviews posted on Tripadvisor on the 20 European airlines with the highest passenger volumes in 2023, according to the CAPA Centre for Aviation [^6] List of Europe’s top 20 airlines: https://centreforaviation.com/analysis/reports/europes-top-20-airline-groups-by-pax-2023-ryanairs-lead-is-set-to-endure-68011. We collected reviews posted between January 1, 2023 and September 24, 2024. This period was selected as it follows the lifting of hygiene measures related to the COVID-19 pandemic, which had been a frequent topic in earlier reviews. A maximum of the first 300 sub-pages listing the reviews on each airline were crawled.
In total, 15,493 reviews were gathered. Non-English reviews were filtered out using langdetect777langdetect: https://pypi.org/project/langdetect, resulting in 15,483 reviews. Named entity recognition (NER) was applied using spaCy (en_core_news_lg model) Honnibal and Montani (2017) to anonymize references to locations, personal names, and time-related information. Identified entities were replaced with placeholders "LOC", "PERSON", and "DATE". Finally, the reviews were segmented into 52,098 sentences using the NLTK Tokenizer Loper and Bird (2002).
B.2 Data Annotation
4,000 sentences were randomly chosen from the 52,098 sentences. We ensure that there is an equal number of sentences from reviews with 1-, 2-, 3-, 4-, or 5-star ratings in order to achieve, that the number of aspects expressing positive, negative, or neutral sentiment is equal to some extent.
We aimed to obtain about 2,000 sentences, acknowledging that some of the 4,000 sentences might (1) not address any of the considered aspect categories, (2) not express any sentiment towards at least one of the considered aspect categories, (3) not be in English but in another language, or (4) be incorrectly tokenized by the NLTK tokenizer, with the annotators identifying multiple sentences instead of one.
B.2.1 Annotation Task
In line with Zhang et al. (2021a), all opinion expressions were annotated in the format of (a, c, o, p)-quadruples. Similar to the ASQP datasets Rest15 and Rest16 introduced by Zhang et al. (2021a), a total of 13 aspect categories were considered for annotation. These are as follows:
AIRLINE#GENERAL
AIRLINE#PRICE
AIRLINE#SERVICE
AIRPORT-OPERATION#BAGGAGE
AIRPORT-OPERATION#BOARDING
AIRPORT-OPERATION#CHECK-IN
ONBOARD#CLEANLINESS
ONBOARD#ENTERTAINMENT
ONBOARD#FOOD
ONBOARD#PRICE
ONBOARD#SEAT-COMFORT
ONBOARD#SEAT-LEGROOM
PUNCTUALITY#GENERAL
Reviewers on Tripadvisor are provided with multiple evaluation criteria and can optionally rate those on a scale of one to five stars in addition to submitting a written review. Since we intended to consider 13 aspect categories, similarly to Zhang et al. (2021a), we adapted Tripadvisor’s nine categories and made several modifications. Specifically, we divided the ’Check-in and boarding’ category into two distinct aspects of a parent category named ’Airport Operation’ and also added an attribute ’baggage’, resulting in three attributes of that parent category. The ’Price’ category was further refined to separately consider the price of onboard offers and the airline’s overall pricing.
Since we capture various aspects related to different aspects of the onboard experience, we did not include a separate ’Onboard Experience’ category, as it can be found on Tripadvisor. Instead, we considered an additional category, PUNCTUALITY#GENERAL. Song et al. (2020) demonstrated that flight delays have a significant impact on overall satisfaction with the flight experience. Lastly, we introduced the AIRLINE#GENERAL category to encompass general aspects associated with the airline.
B.2.2 Data Labelling Process
Annotators were provided with an annotation guideline [^8] Annotation guidelines: https://github.com/anonymization, adapted from the SemEval-2015 guideline (Pontiki et al., 2015), with modifications for the airline domain. Instead of examples from restaurant reviews, examples from airline reviews were provided.
Similar to the approach applied for SemEval-2015 by Pontiki et al. (2015), annotator A annotated all 3,700 sentences, while annotator B reviewed the annotations and, where necessary, proposed a revised annotation. Both Annotators A and B were PhD students with prior experience in annotating datasets for ABSA. The annotation process was conducted using Google Sheets [^9] Google Sheets: https://workspace.google.com/products/sheets.
In 115 out of 4,000 sentences, annotator B suggested a different label than annotator A. Of these 115 proposed revised annotation, 79 were accepted by annotator A. For the other 36 suggested revisions, it was jointly decided that in 25 cases the original annotation by Annotator A would be retained and in nine cases, the annotation by annotator B was chosen. For the remaining two sentences, a consensus was reached on an annotation distinct from their initially proposed labels.
Of the 4,000 annotated examples, 61 were excluded since a sentence-splitting error made by the NLTK tokenizer was identified by the annotators. 1,909 were further excluded as no sentiment was expressed towards the considered aspect categories. 99 sentences were excluded due to an error where either sensitive data was not anonymized or parts of a sentence were anonymized where no anonymization was required. Finally, one non-English sentence was excluded. This resulted in a dataset of 1,930 sentences.
B.3 Dataset Properties
The properties of the FlightABSA dataset (training and test sets) are presented in Table 4. A train-test-validation split (70:20:10) of the entire dataset was applied.
Notably, similar to the SemEval datasets, there is a low representation of neutral opinions and implicit aspects. Class-imbalance can also be observed in the aspect categories. For instance, the category ONBOARD#PRICE appears only 17 times in the overall dataset, while the category AIRLINE#SERVICE occurs 986 times.
Appendix C Prompt
Figure 3: Example of a prompt employed for the ASQP task. The prompt comprises an explanation on the considered sentiment elements, output format and annotated examples in the case of few-shot learning.
Appendix D TASD: Performance Scores
Method Prompting Strategy # Few-Shot / # Train Rest15 Rest16 FlightABSA
OATS
Coursera
OATS
Hotels
F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec
Gemma-2-9B - 0 15.68 17.88 13.99 30.51 35.47 26.80 46.04 50.21 42.80 14.82 16.39 13.52 31.26 38.60 26.30
10 32.21 32.06 32.38 33.82 32.39 35.41 35.63 38.97 32.82 14.67 14.03 15.37 33.87 36.53 31.57
20 37.72 37.56 37.89 40.20 37.59 43.21 37.59 37.76 37.43 17.81 16.68 19.14 32.38 31.85 32.94
30 38.21 38.04 38.39 43.48 41.74 45.38 41.81 40.62 43.06 18.14 16.58 20.04 35.20 34.90 35.52
40 41.01 40.57 41.47 44.90 43.82 46.03 44.62 43.16 46.20 22.16 20.30 24.43 37.25 37.78 36.76
50 41.71 42.06 41.37 46.66 45.50 47.89 45.50 42.76 48.66 19.57 18.01 21.43 39.24 39.90 38.60
SC 0 14.50 18.43 11.95 30.53 38.31 25.38 46.83 55.58 40.45 16.14 20.00 13.52 33.06 45.18 26.07
10 38.75 59.90 28.64 42.13 64.35 31.32 42.61 69.07 30.81 12.07 29.60 7.58 40.17 58.13 30.68
20 47.34 66.31 36.80 51.40 64.78 42.61 46.73 64.98 36.48 18.21 38.93 11.89 43.99 67.66 32.59
30 45.41 62.27 35.74 54.77 70.38 44.82 53.96 73.77 42.53 21.85 42.11 14.75 45.95 64.74 35.61
40 49.03 62.73 40.24 54.02 68.52 44.59 57.24 72.13 47.45 25.21 41.90 18.03 43.56 61.03 33.86
50 49.96 64.96 40.59 54.01 68.21 44.70 57.94 70.16 49.34 22.89 37.91 16.39 47.84 67.54 37.04
Gemma-2-27B - 0 35.98 35.45 36.52 50.89 50.72 51.06 53.88 51.83 56.11 26.84 24.29 30.00 42.45 41.23 43.75
10 51.27 53.37 49.33 55.66 57.77 53.71 55.69 57.29 54.18 31.50 30.45 32.62 48.47 49.72 47.28
20 54.04 55.81 52.38 60.23 61.18 59.32 56.15 55.64 56.67 34.90 33.70 36.19 48.95 49.14 48.78
30 54.71 56.59 52.95 61.14 61.87 60.44 58.27 57.76 58.79 36.38 35.34 37.50 51.84 52.51 51.19
40 55.81 57.49 54.22 61.67 63.14 60.28 60.15 59.16 61.17 37.78 37.18 38.40 52.21 53.50 51.00
50 56.42 58.55 54.44 60.84 62.06 59.67 60.26 58.44 62.19 37.67 37.08 38.28 54.12 55.53 52.78
SC 0 37.08 37.48 36.69 53.13 55.06 51.34 55.29 54.28 56.33 29.03 28.38 29.71 44.34 45.14 43.56
10 55.00 61.37 49.82 61.04 69.02 54.71 58.97 64.85 54.06 35.57 40.74 31.56 50.57 56.47 45.79
20 55.74 61.28 51.12 63.43 68.23 59.25 60.16 63.58 57.09 38.95 43.85 35.04 53.05 57.65 49.13
30 57.44 64.05 52.07 64.18 68.75 60.19 62.59 66.67 58.98 43.62 50.27 38.52 55.27 62.40 49.60
40 60.40 66.76 55.15 66.03 73.13 60.19 64.29 67.43 61.44 42.26 48.41 37.50 55.96 62.90 50.40
50 61.69 68.35 56.21 65.99 72.86 60.30 65.16 67.97 62.57 43.90 50.67 38.73 57.82 64.64 52.31
MVP (Gou et al., 2023) 10 25.08 30.30 21.40 17.08 18.13 16.16 19.04 22.53 16.48 31.83 35.26 29.02 21.16 25.90 17.90
20 32.88 36.45 29.96 28.35 30.55 26.50 23.25 27.35 20.23 34.26 37.99 31.19 27.02 33.95 22.45
30 36.34 40.19 33.18 39.04 41.81 36.62 31.06 35.91 27.37 35.44 38.57 32.79 37.06 42.95 32.59
40 41.07 44.05 38.46 41.04 43.80 38.60 37.20 41.33 33.84 36.76 40.06 33.98 41.56 48.19 36.53
50 42.13 45.54 39.20 44.09 46.91 41.58 44.11 47.92 40.87 37.84 41.35 34.88 44.49 51.12 39.40
800 62.54 64.87 60.38 68.22 69.24 67.24 64.61 64.72 64.50 50.61 51.33 49.92 66.67 67.51 65.85
Full 64.53 - - 72.76 - - 68.67 67.84 69.53 50.97 51.42 50.53 69.37 69.58 69.16
DLO (Hu et al., 2022) 10 8.75 10.72 7.38 6.66 7.61 5.93 8.82 10.44 7.64 15.94 17.69 14.51 14.91 18.74 12.39
20 21.09 21.04 21.40 18.05 16.87 19.53 8.60 10.18 7.45 20.38 22.11 18.93 18.84 21.21 17.20
30 21.83 22.03 21.87 17.46 17.18 18.08 12.08 13.27 11.12 22.45 23.51 21.68 25.18 22.47 29.33
40 31.01 33.70 28.76 28.88 29.96 27.90 26.82 30.07 24.23 30.90 34.18 28.20 36.35 38.86 34.36
50 36.92 39.57 34.60 35.87 37.18 34.66 33.57 36.10 31.38 34.26 37.64 31.43 40.10 45.21 36.05
800 61.54 63.54 59.67 69.31 69.37 69.25 67.69 69.41 66.05 51.36 52.58 50.20 67.48 68.80 66.21
Full 63.06 - - 71.97 - - 69.74 70.22 69.26 51.86 52.73 51.02 67.70 68.41 67.01
Paraphrase (Zhang et al., 2021a) 10 15.92 19.40 13.49 12.78 12.50 13.11 15.94 18.86 13.80 23.51 26.10 21.39 18.05 18.51 17.62
20 25.97 24.10 28.28 24.73 20.81 30.55 22.62 26.77 19.58 29.00 32.12 26.43 26.46 26.03 26.96
30 30.20 29.69 30.75 35.05 33.20 37.16 30.11 32.22 28.28 33.19 36.00 30.78 37.44 38.05 36.85
40 34.98 35.84 34.18 39.01 38.67 39.37 33.30 34.35 32.33 34.42 37.50 31.80 39.98 43.05 37.33
50 39.53 40.47 38.63 41.87 42.31 41.44 39.61 39.54 39.70 35.74 38.98 32.99 43.26 46.88 40.16
800 62.55 64.10 61.07 69.28 69.23 69.34 68.14 68.35 67.94 51.14 51.63 50.66 68.29 68.57 68.01
Full 62.95 - - 71.79 - - 69.03 68.72 69.34 52.34 52.56 52.13 69.11 68.82 69.41
Table 5: Performance scores for TASD. For the Rest15 and Rest16 datasets, performance scores achieved when employing the full training set ("Full") are taken from Gou et al. (2023), Hu et al. (2022) and Zhang et al. (2021b) for MVP, DLO and Paraphrase, respectively. The best score achieved by a method is presented in bold.
Appendix E Element-Level Performance Scores
E.1 TASD
Sentiment Element Prompting Strategy # Few-Shot / # Train Rest15 Rest16 FlightABSA
OATS
Coursera
OATS
Hotels
F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec
Aspect Term - 0 63.26 53.03 78.38 69.02 61.16 79.22 63.97 54.61 77.23 65.00 57.62 74.55 64.01 54.29 77.99
10 70.00 66.15 74.35 73.68 72.05 75.52 74.41 70.16 79.21 65.74 65.68 65.83 67.89 63.54 72.91
20 71.37 66.72 76.72 75.72 74.06 77.48 75.21 68.27 83.72 67.70 64.47 71.47 69.46 63.11 77.25
30 71.15 67.40 75.35 76.32 73.72 79.15 74.91 68.26 83.01 69.64 65.53 74.31 70.35 65.30 76.25
40 71.07 68.95 73.32 75.86 76.36 75.39 74.46 68.77 81.19 70.62 66.04 75.88 70.49 67.11 74.23
50 70.59 68.54 72.77 75.63 75.35 75.95 74.93 70.01 80.59 70.19 64.62 76.82 71.62 68.22 75.38
SC 0 64.98 55.66 78.04 71.23 65.35 78.27 64.99 57.78 74.26 66.38 60.08 74.17 65.48 57.79 75.51
10 72.63 71.84 73.43 76.21 79.09 73.53 75.54 74.66 76.44 67.82 70.98 64.93 67.88 67.57 68.19
20 72.43 70.77 74.17 78.66 80.62 76.80 75.49 71.30 80.20 69.24 68.36 70.14 72.32 70.59 74.14
30 72.29 71.90 72.69 78.68 79.27 78.10 76.05 71.83 80.79 73.60 74.22 72.99 71.13 73.01 69.34
40 73.63 75.19 72.14 76.42 82.42 71.24 76.38 73.58 79.41 73.47 72.79 74.17 72.62 74.64 70.71
50 73.26 74.81 71.77 76.45 80.99 72.39 76.82 75.29 78.42 74.07 72.40 75.83 71.93 74.21 69.79
Aspect Category - 0 60.30 60.60 60.00 76.05 76.08 76.02 67.04 66.06 68.06 77.11 75.60 78.68 65.77 65.13 66.44
10 71.67 74.95 68.67 76.90 78.56 75.32 75.27 75.81 74.74 75.41 77.73 73.24 66.43 68.89 64.14
20 73.72 76.99 70.71 80.45 81.08 79.84 76.92 77.78 76.09 77.25 76.67 77.84 68.02 69.00 67.08
30 75.24 78.36 72.36 81.84 82.62 81.08 78.97 79.81 78.16 79.84 79.81 79.88 71.19 73.10 69.38
40 76.86 80.05 73.91 81.44 82.85 80.08 80.29 80.86 79.74 80.63 80.03 81.24 71.86 74.49 69.42
50 77.15 80.82 73.81 80.69 81.70 79.70 79.83 80.46 79.21 80.58 79.14 82.08 71.69 74.12 69.42
SC 0 60.17 61.85 58.58 75.52 77.67 73.49 67.22 68.97 65.55 76.69 76.39 77.00 66.09 68.58 63.77
10 72.43 81.12 65.42 78.27 86.11 71.74 76.67 81.11 72.70 75.35 82.42 69.40 67.53 75.46 61.10
20 73.71 81.76 67.10 81.63 87.04 76.85 78.61 83.44 74.31 77.68 82.00 73.80 69.68 76.09 64.27
30 74.93 84.22 67.48 82.38 88.06 77.39 80.55 85.06 76.50 80.93 86.04 76.40 73.15 83.19 65.28
40 78.15 87.14 70.84 82.76 90.03 76.58 82.04 86.93 77.66 81.39 85.81 77.40 72.98 82.22 65.61
50 78.68 88.67 70.71 82.52 89.47 76.58 81.85 86.89 77.37 81.85 86.09 78.00 73.87 82.64 66.78
Sentiment Polarity - 0 89.49 90.01 88.98 90.34 89.58 91.11 93.29 93.97 92.62 91.24 90.60 91.88 90.47 90.48 90.47
10 90.07 91.91 88.30 89.81 91.48 88.19 94.05 95.24 92.89 89.69 91.81 87.68 88.26 92.06 84.76
20 90.95 92.32 89.63 92.19 92.96 91.43 94.19 95.57 92.85 91.83 91.92 91.74 91.59 92.21 90.99
30 91.23 92.27 90.22 92.53 93.26 91.81 94.63 95.90 93.40 92.43 92.05 92.81 91.79 93.15 90.47
40 91.13 92.45 89.85 92.62 93.47 91.78 94.43 95.21 93.68 93.06 92.72 93.40 91.54 92.53 90.57
50 90.51 92.17 88.92 92.82 93.53 92.13 94.93 95.79 94.09 92.86 91.66 94.08 92.94 94.36 91.56
SC 0 88.54 91.15 86.07 88.78 90.31 87.30 91.36 95.35 87.69 90.26 91.04 89.49 88.37 91.86 85.14
10 87.07 94.39 80.80 86.94 94.76 80.32 91.45 96.58 86.84 87.61 93.85 82.15 85.49 94.83 77.83
20 87.53 93.81 82.04 89.56 94.53 85.08 91.46 96.40 87.01 89.32 94.29 84.84 89.28 94.71 84.43
30 88.25 94.84 82.51 89.06 94.00 84.60 93.14 97.21 89.40 90.70 94.68 87.04 87.56 94.78 81.37
40 88.03 94.34 82.51 89.54 95.50 84.29 92.60 96.83 88.72 90.56 94.67 86.80 89.53 94.97 84.67
50 87.40 94.11 81.58 89.90 95.70 84.76 93.13 97.39 89.23 90.59 94.43 87.04 89.81 96.23 84.20
Table 6: Gemma-2-27B: Performance scores at element-level for the TASD task. The best score achieved by a method is presented in bold.
Sentiment Element Prompting Strategy # Few-Shot / # Train Rest15 Rest16 FlightABSA
OATS
Coursera
OATS
Hotels
F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec
Aspect Term - 0 51.80 47.39 57.23 54.53 53.23 55.98 49.94 45.89 54.77 57.86 56.47 59.72 52.79 53.45 52.22
10 56.98 53.66 61.00 61.41 58.23 65.07 65.75 64.94 66.61 49.48 56.99 43.79 52.03 57.22 47.78
20 61.04 57.29 65.39 63.21 61.08 65.59 67.36 61.36 74.69 53.83 55.18 52.56 54.10 50.46 58.40
30 63.05 59.72 66.79 66.51 65.71 67.35 67.95 61.06 76.63 56.92 55.01 59.00 56.37 54.36 58.58
40 63.97 62.27 65.79 67.00 65.76 68.30 67.63 63.55 72.40 60.03 57.01 63.46 56.64 55.27 58.12
50 63.68 61.94 65.57 67.42 64.96 70.16 68.13 64.28 72.51 57.83 52.95 63.79 58.59 57.15 60.14
SC 0 51.70 51.56 51.85 53.44 55.56 51.47 52.05 51.25 52.87 58.81 61.72 56.16 55.25 60.88 50.57
10 56.07 76.43 44.28 61.19 82.32 48.69 64.20 80.78 53.27 48.51 79.89 34.83 49.93 71.19 38.44
20 67.25 83.15 56.46 67.58 83.61 56.70 69.49 74.72 64.95 56.04 75.50 44.55 56.51 79.92 43.71
30 66.16 81.40 55.72 69.41 86.76 57.84 71.26 75.44 67.52 60.50 79.54 48.82 58.52 77.15 47.14
40 67.67 81.35 57.93 69.47 83.83 59.31 70.65 78.31 64.36 66.21 78.76 57.11 57.14 76.05 45.77
50 67.94 79.65 59.23 70.72 82.71 61.76 71.00 76.91 65.94 64.62 74.01 57.35 59.55 77.09 48.51
Aspect Category - 0 41.03 46.58 36.70 60.47 67.13 55.05 46.14 50.82 42.25 73.39 79.74 68.20 53.53 65.45 45.31
10 55.53 54.03 57.16 55.70 51.49 60.70 64.72 64.20 65.26 59.92 64.48 55.96 54.34 58.22 50.95
20 60.05 59.24 60.90 62.61 56.89 69.64 70.26 69.55 71.01 64.95 64.21 65.72 50.65 49.57 51.79
30 60.89 59.77 62.06 67.93 63.33 73.24 68.58 65.25 72.29 69.49 67.66 71.44 56.29 55.72 56.89
40 64.12 62.83 65.47 66.81 63.73 70.20 73.01 70.19 76.09 70.68 68.22 73.36 58.69 59.15 58.26
50 63.83 63.94 63.74 67.86 65.03 70.96 73.14 70.69 75.77 70.80 67.31 74.72 59.14 59.68 58.63
SC 0 37.11 45.99 31.10 59.72 69.98 52.09 43.55 51.08 37.96 72.81 84.43 64.00 52.66 70.39 42.07
10 53.69 80.05 40.39 54.24 77.50 41.72 66.73 85.78 54.60 56.56 89.22 41.40 58.53 82.87 45.24
20 60.61 82.96 47.74 64.56 78.05 55.05 72.25 85.92 62.34 63.38 85.96 50.20 55.01 82.61 41.24
30 58.90 78.96 46.97 68.70 85.09 57.60 71.43 82.85 62.77 66.00 88.55 52.60 60.17 83.88 46.91
40 63.08 79.84 52.13 66.29 81.82 55.72 75.35 87.31 66.28 70.11 87.69 58.40 59.30 81.58 46.58
50 65.45 84.49 53.42 67.31 82.87 56.66 75.54 84.48 68.32 69.95 84.66 59.60 62.05 85.84 48.58
Sentiment Polarity - 0 85.30 90.82 80.43 86.10 90.06 82.48 90.97 94.98 87.28 89.32 91.27 87.48 89.48 93.67 85.66
10 88.31 89.87 86.81 87.18 87.35 87.02 88.58 89.41 87.76 86.28 89.06 83.67 85.94 91.99 80.66
20 88.52 89.94 87.15 88.17 86.72 89.68 89.16 89.34 88.99 84.35 83.31 85.43 90.69 91.85 89.58
30 87.70 88.12 87.28 88.05 86.55 89.62 91.31 91.11 91.52 84.61 83.27 86.01 90.67 92.29 89.10
40 88.82 89.54 88.11 88.77 87.51 90.06 89.92 89.21 90.67 87.27 85.80 88.80 90.97 92.12 89.86
50 87.81 89.33 86.35 89.86 88.74 91.02 90.50 89.71 91.32 87.87 85.66 90.22 91.46 92.78 90.19
SC 0 81.69 91.88 73.53 84.54 91.50 78.57 87.74 95.20 81.37 87.92 92.68 83.62 85.60 95.10 77.83
10 68.00 96.58 52.48 67.77 96.48 52.22 76.92 96.15 64.10 67.19 95.50 51.83 76.79 97.81 63.21
20 72.76 94.55 59.13 77.29 95.13 65.08 83.41 96.41 73.50 74.85 94.76 61.86 74.45 97.70 60.14
30 74.10 96.29 60.22 78.12 95.63 66.03 84.47 96.90 74.87 75.52 96.93 61.86 78.09 96.53 65.57
40 78.31 96.38 65.94 78.39 95.44 66.51 84.50 96.70 75.04 79.20 95.83 67.48 79.05 96.92 66.75
50 76.98 95.00 64.71 79.33 96.58 67.30 87.03 96.66 79.15 83.08 97.06 72.62 79.72 97.94 67.22
Table 7: Gemma-2-9B: Performance scores at element-level for the TASD task. The best score achieved by a method is presented in bold.
E.2 ASQP
Sentiment Element Prompting Strategy # Few-Shot / # Train Rest15 Rest16 FlightABSA
OATS
Coursera
OATS
Hotels
F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec
Aspect Term - 0 65.60 57.02 77.23 71.30 64.43 79.82 68.35 59.08 81.07 64.69 59.45 70.95 66.79 59.18 76.66
10 67.38 60.12 76.64 74.53 67.96 82.52 73.29 67.71 79.88 66.65 63.79 69.81 67.75 63.78 72.27
20 68.59 63.37 74.78 74.87 68.39 82.73 73.67 66.89 81.98 67.44 63.48 71.94 68.24 62.44 75.24
30 67.94 62.89 73.87 75.46 69.60 82.41 74.12 67.40 82.34 69.54 64.86 74.98 69.65 64.41 75.84
40 68.27 63.97 73.20 76.19 71.46 81.60 74.93 70.10 80.48 69.32 64.84 74.45 67.61 63.59 72.17
50 69.42 64.93 74.58 75.93 72.80 79.36 74.43 70.89 78.34 69.57 64.89 74.98 68.17 66.55 69.89
SC 0 67.57 62.23 73.91 74.19 71.17 77.48 68.73 63.68 74.65 66.97 64.68 69.43 68.78 63.80 74.60
10 70.68 68.39 73.12 75.94 76.34 75.53 71.63 72.80 70.50 67.33 71.05 63.98 67.81 72.89 63.39
20 69.35 70.55 68.18 77.37 76.56 78.19 75.48 76.32 74.65 69.48 72.92 66.35 69.68 74.80 65.22
30 69.24 71.19 67.39 76.62 76.83 76.42 74.08 74.30 73.86 71.27 74.29 68.48 67.72 72.09 63.84
40 70.15 72.92 67.59 77.76 79.34 76.24 74.38 77.75 71.29 71.36 74.48 68.48 66.83 74.09 60.87
50 70.53 75.45 66.21 75.66 80.85 71.10 74.97 80.32 70.30 72.07 76.05 68.48 66.14 78.37 57.21
Opinion Term - 0 62.52 60.46 64.73 65.57 64.22 66.98 51.88 53.34 50.51 57.47 58.07 56.90 44.13 46.48 42.01
10 62.29 58.89 66.10 67.53 64.50 70.87 58.66 58.71 58.60 58.60 58.13 59.08 47.95 50.52 45.64
20 56.77 55.03 58.63 68.57 66.23 71.09 58.04 58.81 57.29 58.06 57.30 58.84 50.20 51.05 49.39
30 59.05 57.00 61.26 67.86 65.79 70.08 54.89 54.80 54.98 61.24 60.46 62.05 53.09 54.07 52.15
40 59.42 56.49 62.69 67.96 66.31 69.70 57.34 57.56 57.13 59.66 58.87 60.49 51.42 52.35 50.52
50 59.74 57.19 62.53 67.93 66.48 69.46 55.79 56.27 55.31 62.21 60.84 63.63 52.02 54.29 49.94
SC 0 62.76 64.45 61.16 65.73 68.18 63.45 52.10 58.10 47.22 58.03 62.05 54.50 44.69 50.00 40.41
10 63.51 67.01 60.35 67.05 70.91 63.59 59.44 66.67 53.62 59.41 66.89 53.44 48.32 59.45 40.70
20 55.66 61.94 50.54 67.54 72.36 63.32 58.08 67.09 51.21 58.01 65.56 52.03 52.50 64.35 44.33
30 60.01 66.67 54.57 67.40 72.76 62.77 55.90 63.78 49.76 61.90 70.75 55.03 53.40 64.34 45.64
40 60.41 67.05 54.97 67.78 72.93 63.32 58.36 67.84 51.21 61.67 69.87 55.20 52.42 64.74 44.04
50 58.05 65.87 51.88 67.21 74.22 61.41 55.73 65.63 48.43 61.80 70.18 55.20 52.63 68.13 42.88
Aspect Category - 0 59.61 59.24 60.00 66.46 66.11 66.82 66.28 65.02 67.59 74.33 73.72 74.96 64.63 65.06 64.21
10 72.31 71.55 73.09 70.99 69.51 72.53 71.24 69.52 73.05 76.01 75.11 76.96 64.58 66.58 62.70
20 72.41 73.00 71.83 77.21 76.69 77.75 75.48 75.57 75.39 77.30 75.31 79.40 67.07 67.89 66.28
30 74.19 74.79 73.61 79.82 79.81 79.82 78.68 77.98 79.39 79.18 77.78 80.64 70.20 71.47 68.98
40 75.51 75.37 75.64 80.86 81.21 80.53 79.24 78.75 79.74 79.20 77.56 80.92 70.32 71.78 68.91
50 75.54 75.82 75.27 81.03 81.66 80.41 78.93 78.78 79.09 78.87 77.07 80.76 69.88 73.07 66.98
SC 0 57.86 61.55 54.58 64.35 68.03 61.05 65.33 68.99 62.04 72.63 75.70 69.80 64.96 69.11 61.27
10 71.23 78.92 64.90 69.91 75.42 65.15 70.77 77.11 65.40 72.99 81.33 66.20 63.84 75.57 55.26
20 69.93 81.37 61.32 78.05 85.79 71.60 74.92 84.71 67.15 75.00 82.14 69.00 66.41 79.35 57.10
30 73.41 85.23 64.47 79.94 88.89 72.62 77.87 87.16 70.36 76.16 84.98 69.00 67.12 80.09 57.76
40 72.98 84.82 64.04 80.64 89.32 73.50 77.90 89.10 69.20 76.81 84.78 70.20 67.85 82.89 57.43
50 73.53 88.51 62.89 79.67 90.04 71.45 77.00 88.45 68.18 76.63 85.86 69.20 66.94 84.87 55.26
Sentiment Polarity - 0 90.19 90.02 90.36 90.99 90.54 91.45 92.75 93.05 92.44 89.52 90.30 88.75 90.13 90.85 89.43
10 90.41 89.89 90.94 91.68 90.86 92.52 93.78 94.27 93.30 90.45 91.49 89.44 86.88 91.19 82.97
20 90.34 89.74 90.94 92.58 92.71 92.45 94.23 95.28 93.20 91.79 91.46 92.13 89.50 91.40 87.69
30 90.79 90.78 90.80 93.23 93.42 93.03 94.88 95.63 94.15 93.02 92.51 93.55 90.33 93.14 87.69
40 91.56 91.16 91.97 94.10 94.17 94.03 94.34 94.73 93.95 92.88 92.03 93.74 89.97 92.27 87.78
50 91.38 91.24 91.52 93.41 93.69 93.14 94.44 94.86 94.02 93.28 92.54 94.03 90.84 94.36 87.59
SC 0 86.92 91.34 82.91 88.75 92.84 85.00 87.75 94.30 82.05 87.26 92.12 82.89 87.95 92.91 83.49
10 85.66 92.64 79.66 86.62 93.95 80.34 87.92 96.33 80.85 84.64 94.29 76.77 82.38 93.96 73.35
20 83.14 93.21 75.04 87.02 94.91 80.34 87.25 97.47 78.97 85.03 93.81 77.75 81.50 94.41 71.70
30 84.04 93.88 76.07 88.01 96.31 81.03 87.09 97.06 78.97 84.71 94.85 76.53 81.63 96.46 70.75
40 83.62 94.41 75.04 87.62 96.09 80.52 85.47 96.96 76.41 86.13 94.72 78.97 81.78 95.58 71.46
50 82.01 94.43 72.48 85.96 95.57 78.10 86.82 97.45 78.29 85.79 96.06 77.51 80.72 97.02 69.10
Table 8: Gemma-2-27B: Performance scores at element-level for the ASQP task. The best score achieved by a method is presented in bold.
Sentiment Element Prompting Strategy # Few-Shot / # Train Rest15 Rest16 FlightABSA
OATS
Coursera
OATS
Hotels
F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec F1 Pre Rec
Aspect Term - 0 57.66 51.95 64.82 63.37 59.28 68.09 55.61 50.85 61.43 60.76 56.90 65.21 55.85 52.81 59.31
10 53.31 48.11 59.88 60.21 56.86 64.01 63.76 63.79 63.80 53.05 55.76 50.85 53.74 60.24 48.56
20 58.40 55.33 61.86 62.57 60.07 65.32 65.63 64.08 67.29 53.20 54.33 52.13 55.80 55.87 55.79
30 59.50 56.25 63.20 65.79 63.04 68.79 66.31 63.12 69.86 57.19 57.06 57.39 57.97 59.66 56.38
40 58.89 57.23 60.67 67.40 66.97 67.84 67.05 66.60 67.52 59.35 57.90 60.95 57.96 59.63 56.38
50 61.39 60.27 62.57 68.40 68.15 68.65 68.71 69.57 67.88 60.17 56.92 63.84 59.99 62.49 57.71
SC 0 60.56 59.13 62.06 63.87 65.79 62.06 55.24 58.08 52.67 63.56 64.10 63.03 56.57 58.58 54.69
10 48.92 77.35 35.77 48.46 87.50 33.51 55.05 81.40 41.58 43.63 82.78 29.62 42.16 80.13 28.60
20 54.64 83.06 40.71 57.99 87.19 43.44 62.83 85.37 49.70 45.81 82.21 31.75 45.89 91.16 30.66
30 57.55 84.35 43.68 62.61 85.80 49.29 62.50 84.75 49.50 56.14 85.85 41.71 50.25 88.95 35.01
40 59.49 84.67 45.85 62.73 87.34 48.94 61.08 87.45 46.93 58.54 82.05 45.50 52.65 88.17 37.53
50 59.69 86.19 45.65 63.86 88.92 49.82 62.86 87.32 49.11 59.18 82.28 46.21 54.31 89.95 38.90
Opinion Term - 0 52.53 55.87 49.57 57.27 60.43 54.43 43.26 49.94 38.16 53.28 57.82 49.42 39.94 47.49 34.48
10 50.58 48.62 52.72 54.93 50.87 59.70 47.79 48.72 46.91 48.71 49.70 47.76 37.17 43.52 32.44
20 50.39 49.31 51.53 56.71 54.44 59.18 48.58 50.44 46.86 50.90 50.20 51.64 42.90 44.81 41.16
30 50.30 48.96 51.72 59.83 58.21 61.55 47.25 48.42 46.16 55.15 54.85 55.45 43.97 47.00 41.31
40 52.35 50.60 54.25 60.10 58.73 61.55 48.45 49.56 47.42 54.79 54.97 54.64 45.45 49.28 42.18
50 54.40 53.27 55.59 60.31 58.75 61.98 49.04 50.25 47.90 57.14 56.73 57.57 47.89 52.46 44.07
SC 0 52.45 60.82 46.10 56.61 66.18 49.46 41.34 55.74 32.85 55.03 64.83 47.80 40.11 52.98 32.27
10 40.31 75.09 27.55 39.92 75.19 27.17 41.32 69.12 29.47 38.56 78.38 25.57 31.01 65.12 20.35
20 41.98 75.79 29.03 48.92 79.51 35.33 45.88 70.78 33.94 40.99 77.45 27.87 33.59 71.56 21.95
30 46.55 75.76 33.60 53.81 79.16 40.76 44.94 70.54 32.97 49.33 79.30 35.80 39.23 74.69 26.60
40 47.06 74.42 34.41 53.01 78.25 40.08 45.06 73.50 32.49 49.76 76.56 36.86 39.41 72.66 27.03
50 46.41 75.68 33.47 54.17 78.21 41.44 47.65 72.41 35.51 50.77 76.79 37.92 42.01 74.72 29.22
Aspect Category - 0 41.78 45.12 38.91 44.29 46.54 42.25 43.52 47.05 40.50 74.10 77.42 71.08 54.21 61.08 48.75
10 56.62 53.15 60.60 53.13 48.89 58.18 60.75 60.04 61.49 60.97 60.44 61.52 50.45 56.73 45.44
20 60.58 58.76 62.55 58.58 56.02 61.41 68.74 70.29 67.27 64.15 61.38 67.20 53.90 55.40 52.49
30 62.57 61.15 64.07 64.22 62.38 66.18 66.60 65.62 67.62 69.59 67.88 71.40 57.09 59.29 55.06
40 64.31 62.65 66.07 61.71 59.92 63.63 69.79 68.52 71.12 71.64 69.67 73.76 58.36 61.67 55.39
50 64.16 62.96 65.42 65.36 64.40 66.35 71.17 70.22 72.15 71.12 68.66 73.76 58.74 62.20 55.66
SC 0 39.11 45.86 34.10 40.17 46.62 35.29 38.58 47.64 32.41 73.64 82.96 66.20 50.89 62.90 42.74
10 47.45 86.69 32.66 40.89 75.00 28.11 56.35 87.93 41.46 45.60 89.47 30.60 44.28 89.80 29.38
20 50.87 89.53 35.53 49.95 81.46 36.02 61.57 89.20 47.01 50.36 89.74 35.00 45.06 93.19 29.72
30 52.39 85.95 37.68 57.14 84.97 43.05 61.27 90.08 46.42 58.34 90.72 43.00 49.70 90.71 34.22
40 57.14 90.12 41.83 54.11 79.55 41.00 61.85 93.24 46.28 61.38 90.62 46.40 49.64 87.03 34.72
50 54.89 88.25 39.83 57.88 84.31 44.07 64.28 90.69 49.78 63.01 90.94 48.20 51.94 87.70 36.89
Sentiment Polarity - 0 86.90 90.61 83.49 88.05 90.90 85.38 90.23 94.07 86.70 88.68 90.66 86.80 89.57 92.39 86.93
10 84.13 83.12 85.16 84.86 81.11 88.97 86.71 87.49 85.95 85.78 85.59 85.97 81.61 91.41 73.73
20 85.33 82.93 87.90 88.25 86.84 89.72 88.80 90.37 87.28 82.77 79.08 86.85 87.34 91.57 83.49
30 86.98 85.83 88.17 88.12 86.80 89.48 90.04 91.41 88.72 84.05 81.43 86.85 88.52 92.31 85.05
40 87.46 86.57 88.38 88.94 88.33 89.55 87.97 87.47 88.48 87.34 85.57 89.19 88.78 92.99 84.95
50 87.68 87.27 88.10 89.50 88.94 90.07 88.69 88.42 88.96 87.65 85.50 89.93 89.00 93.46 84.95
SC 0 83.29 93.04 75.38 83.35 92.99 75.52 82.75 95.53 72.99 85.45 92.09 79.71 85.23 94.54 77.59
10 56.48 97.08 39.83 53.68 97.29 37.07 64.00 96.55 47.86 53.52 95.60 37.16 56.86 97.70 40.09
20 59.88 97.31 43.25 61.81 97.05 45.34 71.30 97.91 56.07 59.08 96.67 42.54 55.16 97.60 38.44
30 61.95 96.39 45.64 67.19 96.45 51.55 68.94 96.90 53.50 64.72 95.69 48.90 61.61 97.45 45.05
40 65.15 97.61 48.89 65.46 94.77 50.00 68.74 98.41 52.82 68.13 95.18 53.06 64.56 98.08 48.11
50 63.22 96.49 47.01 68.88 96.28 53.62 72.47 97.68 57.61 70.09 96.57 55.01 67.39 97.76 51.42
Table 9: Gemma-2-9B: Performance scores at element-level for the ASQP task. The best score achieved by a method is presented in bold.
Appendix F Performance Scores: Visualization
Figure 4: Influence of the amount of few-shot examples on the performance of Gemma-2-9B and Gemma-2-27B. Visualization includes comparison with performance scores of SOTA supervised methods MVP, Paraphrase and DLO.